OpenStudy (anonymous):

factor out the greatest common factor. 2x^2y+6x^2y^2-16xy

OpenStudy (cwrw238):

first find GCF of 2, 6 and 16 can you do that?

OpenStudy (anonymous):

it's 2 right?

OpenStudy (anonymous):

ac

OpenStudy (anonymous):

no 6

OpenStudy (cwrw238):

no it 2 - the greatest number that will divide into 2 6 and 16

OpenStudy (anonymous):

which is 2 right?

OpenStudy (cwrw238):

now find GCF for terms in x and terms in y GCF of x^2, x^2 and x and GCF of y and y^2

OpenStudy (cwrw238):

2 is right yes

OpenStudy (anonymous):

that is where i get confused at

OpenStudy (cwrw238):

GCF of x^2 and x is easy write them as x*x and x - it must be x right?

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

x * x and x - the factor common to both is x

OpenStudy (cwrw238):

same with y^2 and y - gotta be y

OpenStudy (cwrw238):

so GCF we are looking for is 2xy

OpenStudy (anonymous):

2xy is going to be my answer right?

OpenStudy (anonymous):

?

OpenStudy (cwrw238):

no you need to factor the expression now divide 2xy into 2x^2y+6x^2y^2-16xy ill start you off : 2x^2 y / 2xy = x so we get 2xy( x + - ) - try filling in the blanks

OpenStudy (anonymous):

4-2? maybe. im not sure

OpenStudy (anonymous):

i just suck lol

OpenStudy (anonymous):

u still there?

OpenStudy (amtran_bus):

Hi. Listen, have an I can do it attitude! 2 is the largest number that will go into the function so we will factor it out. Is there supposed to be a + or - sign after 2x^2 though?

OpenStudy (anonymous):

nope it's just y then there is a + sign

OpenStudy (amtran_bus):

Oh well. So, what number will go into all of them evenly? Could a variable go into them each also? Tell me.

OpenStudy (cwrw238):

6x^2y^2 / 2xy = 3xy

OpenStudy (cwrw238):

and 16xy / 2xy = 8

OpenStudy (anonymous):

so the answer is 8xy

OpenStudy (anonymous):

?

OpenStudy (cwrw238):

no what goes in the brackets is the result of dividing the whole expression by 2xy and 2xy is outside the brackets

OpenStudy (anonymous):

Ok u lost me

OpenStudy (cwrw238):

2x^2y+6x^2y^2-16xy = 2xy(x + 3xy - 8)

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

this process is called factoring

OpenStudy (anonymous):

so is that the answer or no?

OpenStudy (anonymous):

?

OpenStudy (cwrw238):

yes